Science Report – April 18th

Science Report 18 April 2019
Crew 210
Submitted by Paul Sokoloff

Today we continued to process the soil samples taken at each of the 12 transects we conducted over the last few days. In the Science Dome, we measured the pH and salinity and characterized the texture of the soil samples. These data will be used to characterize the ecotones of the deserts around MDRS in an upcoming paper.

We continued to collect new plant specimens opportunistically, including an annual invasive grass, (Annual Wheatgrass – Eremopyrum triticeum) newly recorded for MDRS.

We plan on continuing our lab work and wrapping up data recording for the plant project tomorrow.

Greenhab Report – April 18th

GreenHab Officer: David Murray

Environmental control: Ambient, fan on from 0930 to 1900; the door closed.

Shade Cloth: 80%

Low Temperature: 24 C

High Temperature: 33 C

Relative humidity: 15%

Hours of Supplemental Light: 0

Daily Water Usage for Crops: 38 L

Water in Blue Tank: Refilling tank in about one hour

Time(s) of Watering: 1030

Changes to Crops: The tomatoes didn’t like the heat today but they’re alright. Tomato starts haven’t shown much growth but everything else is responding very well to the fertilizer.

Narrative: The door was left closed with the fan on high from 0930 to 1900 with passive exhaust as part of an experiment. The internal temperature stayed warm but fluctuated less drastically than it usually does with the door open.

Although the high was above 30 degrees C, I think the average temperature was right around perfect for all of the crops but the tomatoes (and possibly the beans).

Harvested basil, sage, carrot, and parsley.

Harvest: 7 g basil, 7 g parsley, 15 g carrot, 1 g sage

Operations Report – April 18th

Crew 210 Operations Report 18-APR-2019

SOL: 6

Name of person filing report: Shannon Rupert

Non-nominal systems: nothing to report

Notes on non-nominal systems: nothing to report

Generator (hours run last night): 10.5

On (last night): 9:30 pm

Off (this morning): 7 am

Solar— SOC 58 % It fell to 41% this afternoon but then rose again so rather than turn on the generator, I am watching it to see how it does the rest of the evening.

Diesel Reading – 36 %

Propane Reading – 49%

Ethanol Free Gasoline – 4 gallons.

Water (auxiliary tank) – 0 gallons

Water (static tank) – 475 gallons

Auxiliary to Static tank transfer– no

Gallons transferred: No

Water in GreenHab – unknown gallons.

Water (loft) – Static to Loft Pump used – yes

Water Meter: not read

Toilet tank emptied: yes

Deimos rover used: Still non-operational
Hours:

Beginning charge:

Ending charge:

Currently charging:

Sojourner rover used: ASSIGNED TO DIRECTOR, yes

Hours: Did not record

Beginning charge: 100

Ending charge: 62

Currently charging: yes

Spirit rover used: yes

Hours: did not check

Beginning charge: 100

Ending charge: did not record

Currently charging: yes

Opportunity rover used: no

Hours: not recorded

Beginning charge: 100

Ending charge: 94

Currently charging: yes

Curiosity rover used: yes

Hours: did not record

Beginning charge: 100

Ending charge: 64

Currently charging: yes

Notes on rovers: Nothing to report

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3) none

Reason for use: n/a

Oil Added? No

ATV Fuel Used: 0 Gals.

# Hours the ATVs were Used today: 0

Notes on ATVs: nothing to report

HabCar used and why, where? Yes, to town for water

CrewCar used and why, where? Yes, south for live outreach event and then to town to pick things up at post office and have lunch

General notes and comments: We added one of the submersible pump to the transfer tank in the Hab Car but have yet to test it.

Summary of internet: Nothing to report

Summary of suits and radios: Nothing to report

Summary of Hab operations: Nothing to report

Summary of GreenHab operations: David will send a full report today.

Summary of ScienceDome operations: We are analyzing our samples in the lab. We collected 60 new plants this week.

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Nothing to report

Questions, concerns, and requests to Mission Support: Nothing tonight, thank you.

Sol Summary – April 18th

Crew 210 Sol Summary Report 18-APR-2019

Sol: 5

Summary Title: Live from the surface of Mars

Author’s name: Paul Sokoloff

Mission Status: Nominal

Sol Activity Summary: This morning through to early afternoon our crew participated in a Livestream with Discover the Universe, engaging over 150 classrooms live with a presentation and Q and A about Mars, the Mars Society, and MDRS. This presentation will soon be uploaded to YouTube where it will continue on as a legacy outreach component of crew 210

(http://www.discovertheuniverse.ca/workshop/special-webinar-live-from-mars-april2019/).

After a water run into Hanksville, our team continued to process our field samples in the Science Dome and will continue to do so tomorrow. By recording physical characteristics of the soil samples taken from our quadrats, we hope to discover the factors driving plant diversity (or the lack thereof) in the different ecological areas around MDRS.

Look Ahead Plan: We plan on continuing our work in the Science Dome tomorrow and will work on putting together a plan for our multi-year biodiversity inventory.

Anomalies in work: No issues.

Weather: Bright and sunny all day.

Crew Physical Status: Good.

EVA: We presented out Livestream from the area just north of "Zubrin’s Head" on the west side of Cow Dung Read, south of the blind corner on the road.

Reports to be filed: GreenHab Report, Operation Report, Science Report, Daily Photo Report

Support Requested: None.